OpenStudy (anonymous):

What is the limit- lim x->-1 [1/sqrt(1+x)]-1 ------------- x

hartnn (hartnn):

is L'Hopital's rule allowed(or have you been taught) ?

OpenStudy (anonymous):

haven't learned it

OpenStudy (anonymous):

I have heard about it a lot recently but I believe we are covering that next chapter

OpenStudy (anonymous):

algebra to the rescue

hartnn (hartnn):

okk...then lets do by conjugate method,

OpenStudy (anonymous):

i have tried algebraically but i get 1/(sqrtx)+2x+x^2

OpenStudy (anonymous):

which i cant answer because we cant really square root -1

hartnn (hartnn):

\(\large \dfrac{1}{\sqrt{x+1}}-1=\dfrac{1-\sqrt{x+1}}{\sqrt{x+1}}\dfrac{1+\sqrt{x+1}}{1+\sqrt{x+1}}\) have you done this only ?

hartnn (hartnn):

then numeratorr will come out to be 1-(x+1)

OpenStudy (anonymous):

i did that to get 1/(sqrt 1+x)(1+sqrt(1+x)

OpenStudy (anonymous):

sorry i meant x/(sqrt 1+x)(1+(sqrt(1_x))

hartnn (hartnn):

wait, why didn't u just substitute x=-1 ??

OpenStudy (anonymous):

because it would give me a 0 in the denominator which is undefined

OpenStudy (anonymous):

but im not too sure about it being undefined.

hartnn (hartnn):

a/0 is not undefined if a is positive, its +infinity if a is negative, its -infinity

OpenStudy (anonymous):

yes but because the two limits are not equal it is considered undefined or non existant

hartnn (hartnn):

here, you need to find the left and right hand limits, you know what they are ?

OpenStudy (anonymous):

the left side of the limit must equal the right side of the limit and -infinity doesn't equal +infinity

OpenStudy (anonymous):

i know what the left and right sided limits are

hartnn (hartnn):

yes, they come out to be unequal, and this limit actually DNE

OpenStudy (anonymous):

ok well then what about this one? lim x->-2 x^2/(x^2-4)

hartnn (hartnn):

right hand limit = -infinity left hand limit = i*infinity

OpenStudy (anonymous):

i stated because we are approaching from the left we get negative infnity

hartnn (hartnn):

for that left hand limit =+infinity rihght hand limit =-infinity hence that limit also DNE

hartnn (hartnn):

could you think of how ?

OpenStudy (anonymous):

i understand how, im just not sure about stating undefined in this question but i suppose it would work

OpenStudy (anonymous):

so i guess for both answers it is negative infinity right?

hartnn (hartnn):

no...both of these limits 'Does not exist' because for both of these, left hand limit not = right hand limit.

hartnn (hartnn):

negative infinity does not imply that limit DNE

OpenStudy (anonymous):

yes the question is only asking for the left side of the limit and the left side of the limit does exist as negative infinity

hartnn (hartnn):

ohh...but for 1st LHL = i/infinity for 2nd, LHL = +infinity

OpenStudy (anonymous):

i thought for both it would be -infnity as it is asking x->-2 and x->-1

OpenStudy (anonymous):

yes i understood that it is -infinity

hartnn (hartnn):

let me explain the easier one, when x->-2 from left means x is less than -2,but very near to it so, x+2 is NEGATIVE but very near to 0 hence 1/(x+2) is −∞ and there is (x-2) also, which is -4, which is negative! so, overall, its +infinity. got this ?

OpenStudy (anonymous):

i am sorry, i typed the question wrong. For the easier one it is asking for the limit x->-2-

OpenStudy (anonymous):

while for the other one it is simply asking for x->-1. So for the easier one it is asking for the left side of the limit while the harder one is just asking for the limit which we found to be DNE

hartnn (hartnn):

yes, \(-2^-\) means less than -2, but very near to it. so, for 2nd its +infinity and for 1st its DNE because LHL not =RHL

hartnn (hartnn):

you got how +infinity ?

OpenStudy (anonymous):

no i am still lost. I learned that −2 − or the left sided limit would mean negative infninty and the right sided limit would be poisitve infnity

hartnn (hartnn):

i heard you saying you got how 1/(x+2) is -infinity, any doubts there ?

OpenStudy (anonymous):

for this question- [1/sqrt(1+x)]-1 ------------- x I know it is DNE because the questions isn't asking for only one side of the limit

hartnn (hartnn):

ok, so lets do that first, to find that limit, we need to find both LHL and RHL for LHL, we thake, \(\large -1^-\) means x is very near to -1 but less than -1 for RHL, we thake, \(\large -1^+\) means x is very near to -1 but greater than -1 got this much ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

good, so when we find LHL and RHL, they don't come out to be =, and hence the original limit DNE. right ? done with 1st ?

hartnn (hartnn):

or you want to know how we found LHL and RHL ?

OpenStudy (anonymous):

ok that i got now about the second question

hartnn (hartnn):

to find 2nd limit, we again need to find both LHL and RHL for LHL, we thake, \(−2^−\) means x is very near to -2 but less than -2 for RHL, we thake, \(−2^+\) means x is very near to -2 but greater than -2 can you find LHL and RHL on your own ?

OpenStudy (anonymous):

yeah LHL is -infinity while RHL is +infinity

hartnn (hartnn):

how ?

OpenStudy (anonymous):

because you cant algebraically change the question

OpenStudy (anonymous):

the question being x^2/x^2-4 with the lim x->-2

hartnn (hartnn):

yes, the function is x^2/x^2-4 = x^2/(x+2)(x-2) and we need to put x=-2 for LHL, \(x=-2^-\), so x+2 will be negative , right ?

OpenStudy (anonymous):

ohhhh i see what i did wrong now

hartnn (hartnn):

see if you get this : x is less than -2, so x+2 is negative, and near to 0 so, 1/(x+2) is -infinity

hartnn (hartnn):

but we still have x-2 !

OpenStudy (anonymous):

x^2/x^2-4 = x^2/(x+2)(x-2) then we change it too- -2^2/(-2+2)(-2-2) which would give is - 4/-4

OpenStudy (anonymous):

no it wont my bad it would still give use 0 h

OpenStudy (anonymous):

and just in case, the limit isn't asking for -2- it is only -2

hartnn (hartnn):

unless you are clear with this, 1/(x+2) is -infinity lets not proceed....

OpenStudy (anonymous):

that i understood if it was asking for the LSL but its not asking for the LSL

hartnn (hartnn):

even if its only -2, we will need to find both LHL(-2-) and RHL(-2+)

hartnn (hartnn):

yes, so, 1/(x+2) is -infinity but x-2 = -4 is also negative! so, negative from -4 and -infinity combines to give you +infinity.

OpenStudy (anonymous):

yes we do, or we can simply plug in 2 itself and find thenswer

hartnn (hartnn):

you meant -2 ? no....we need to find both LHL and RHL.... if they are unequal, limit DNE, as in this case. if you only plug in -, you'll get +infinity, which is incorrect.

hartnn (hartnn):

**only plug in -2,

OpenStudy (anonymous):

yeah my bad i meant -2 and i see what you mean but we would only need to do that if we had two separate functions. We only have one function here and so we only need to plug in one number

OpenStudy (anonymous):

in other words the left sided limit begins at -2 while the right also begins at -2 so there is no need for two equations.

hartnn (hartnn):

there's only one equation, x^2/(x+2)(x-2) but when we take LHL / RHL, we plug in different values, one less than -2, one greater than -2 thats what they mean, if the curve approaches the same value from left and right , only then limit exist, which is not case here, x^2/(x+2)(x-2) approaches +infinity from left and -infinity from right.

hartnn (hartnn):

**from left of -2 **from right of -2

OpenStudy (anonymous):

and it does approach the same value doesn't it?

OpenStudy (anonymous):

i mean if we plug in -1.99999999999 and -2.00000000001 into x we would get almost the same numbers

hartnn (hartnn):

http://www.wolframalpha.com/input/?i=x%5E2%2F%28x%2B2%29%28x-2%29 see the curve at x=-2 from left it goes UP, to +infinity from right , it goes DOWN to -infinity so, no,it does not approach the same value

hartnn (hartnn):